OpenStudy (anonymous):

Solve the following system of equations for y: 2x+y=12 -2x+3y=-4 a)8 b)2 c)3 d)4

OpenStudy (anonymous):

Add the first equation to the second one, you get \(4y=8\) implies that \(y=2\).

OpenStudy (anonymous):

add first equation to the second: 4y=8 solve for y: y=8/4=2 now, substitute value of y in any given equation: 2x+2=12 2x=10 x=5 so, your answer: (5,2)... Actually, you need ONLY value of y = 2